It was found that 90% of cucumber seeds sown in the soil germinate. Determine the most likely

number of germinated grains if there are 70 grains in the package.


1
Expert's answer
2022-05-16T15:33:42-0400

By the binomial distribution, the probability that k seeds germinate is "\begin{pmatrix}
 70 \\
 k
\end{pmatrix}(0.9)^k(1-0.9)^{70-k}"

The expected value of the number of seeds to germinate is "pn=70*0.9=63".

So, we need to compute the binomial distribution probability values for k=62, k=63 and k=64 to find the maximum.

For "k=62"

"P=\frac {70!} {(70-62)!62!} (0.9)^{62}(0.1)^{70-62}=0.1374"

For "k=63"

"P=\frac {70!} {(70-63)!63!} (0.9)^{63}(0.1)^{70-63}=0.15704"

For "k=64"

"P=\frac {70!} {(70-64)!64!} (0.9)^{64}(0.1)^{70-64}=0.15459"

As we can see, "P(k=63)=0.15704" is the most probable.

So, the most likely number of germinated grains if there are 70 grains in the package is 63.